OpenStudy (anonymous):

From least to greatest 3 39/40, 3 19/20, 3 1/2

OpenStudy (anonymous):

3 1/2, 3 19/20, 3 39/40

OpenStudy (anonymous):

\[\left\{\frac{7}{2},\frac{79}{20},\frac{159}{40}\right\} \]